Can anyone help me solve this problem, its says find the solution set for
x2-3x-10 less than or equal to 0

Answer by alanc(27)

The solution set is a set of x-values that satisfy x^2 - 3x - 10 <= 0. It will be an interval.
Factor x^2 - 3x - 10 into (x - 5)*(x + 2), because 2 * (-5) = -10 and 2 - 5 = -3.
First find x values such that (x - 5)*(x + 2) <= 0.
for (x - 5) we see for x <= 5 (x - 5) <= 0
for (x - 5) we see for x > 5 (x - 5) > 0
for (x + 2) we see for x <= -2 (x + 2) <= 0
for (x + 2) we see for x > -2 (x + 2) > 0

draw a line -----------------0++++++++++++++++++++
5
--------0+++++++++++++++++++++++++++++
-2
(x-5)*(x+2) : ++++++++0--------0++++++++++++++++++++



so (x-5)*(x+2) <= 0 when -2 <= x <= 5
Solution set : -2 <= x <=5